OpenStudy (anonymous):

The curve y = ax2 + bx + c touches the line y = x in the origin. It also touches the line y = 2x +3. Determine the function

OpenStudy (anonymous):

by touhes theyy mean tangent to

OpenStudy (anonymous):

i am getting a=-1/12, b=1, c=0, but let me check again

OpenStudy (anonymous):

can u tell me how ?

OpenStudy (anonymous):

well, the tangent touches the curve at origin. so, the curve must go through origin, put x and y as 0, u will get c=0

OpenStudy (anonymous):

yeah i got c=0 and b=1 but im stuck at a

OpenStudy (anonymous):

put y=2x+3 then u will get a quadratic make the discriminant 0 put c=0 and b=1 in that expression.. u will get a

OpenStudy (anonymous):

got it? :)

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

eh! why? u r facing problem in which point?

OpenStudy (anonymous):

what is the discriminant ?

OpenStudy (anonymous):

b^2-4ac

OpenStudy (anonymous):

yeah i get 1+12a

OpenStudy (anonymous):

when the quadratic is: ax^2+by+c=0

OpenStudy (anonymous):

but why is 1+12a=0

OpenStudy (anonymous):

u mean?

OpenStudy (anonymous):

D=12a+1... u said i should make it equal to 0

OpenStudy (anonymous):

means?

OpenStudy (anonymous):

why should i make it =0 ??

OpenStudy (anonymous):

as u have to make the discriminant 0 :)

OpenStudy (anonymous):

why ?what is the rule ???

OpenStudy (anonymous):

substitute y with with the expression of the tangent.. as the two roots will be same, discriminant will be 0. if u dont get it, go through quadratic chapter properly..

OpenStudy (mertsj):

If b = 1 and c = 0 then you have y = ax^2+1x +0 But y = 2x+3 So: 2x+3 = ax^2 +x or ax^2 -x-3 = 0 But b^2-4ac must be 0 so (1)^2 - 4(a)(-3)=0 1+12a = 0 and a = -1/12

OpenStudy (anonymous):

do we make it zero to only get one root ?

OpenStudy (amistre64):

i envision something that looks like this 

OpenStudy (amistre64):

y = ax^2 + bx + c; (0,0), c=0 y = ax^2 + bx y' = 2ax + b; (0,1), b=1 ..................................... y = ax^2 + x; y = 2x+3 ax^2 + x = 2x+3 ax^2 - x - 3 = 0 and since it can only touch in one spot, the determinate must equal 0 1-4(-3)(a)=0 1+12a=0 a = -1/12 i agree .....

OpenStudy (anonymous):

ok now i got it ty all sry i was so annoying

OpenStudy (amistre64):

'sok :)